# Flowing from Words to Pixels: A Framework for Cross-Modality Evolution

### v1

@@ 1 Introduction
In this work, we present key architecture design solutions for overcoming these challenges: First, we employ a Variational Encoder for encoding the source modality data distribution to the same shape as the target modality, and show that the resulting regularization in the source distribution is essential for generation performance. Secondly, we enable CFG in cross-modal flow matching through the introduction of a binary conditioning indicator during training, and demonstrate the quantative benefits of this approach compared to alternative CFG methods. We present CrossFlow; a general framework for mapping between two different modalities without the need for any conditioning mechanism or noise distribution. Typically, different cross-modal generation tasks require task-specific architectural and training modifications, but CrossFlow works for different tasks without any such changes.
Using the ubiquitous albeit challenging text-to-image (T2I) generation task as our primary setting, we show the significant result that CrossFlow outperforms commonly used flow matching baselines, given the same training data, model size, and training budget, all without requiring any cross-attention layers. CrossFlow exhibits improved scaling behavior over standard flow matching using cross-attention when scaling training steps or model size, and is also compatible with a variety of Large Language Models (LLMs), including CLIP [66], T5 [67], and Llama3 [19]. Additionally we demonstrate that since our approach encodes the source distribution into a regularized continuous space with semantic structure, CrossFlow enables exciting new latent arithmetic for the text-to-image task, e.g., $\mathcal{L}$(“A dog with a hat”) + $\mathcal{L}$(“Sunglasses”) – $\mathcal{L}$(“A hat”) creates an image of a dog wearing sunglasses without a hat.
We demonstrate the general-purpose nature of CrossFlow on various cross-modal/intra-modal tasks: image-to-text (image captioning), image-to-depth (depth estimation), and low-resolution to high-resolution image (super-resolution). we show that our CrossFlow approach achieves comparable or even superior performance to various state-of-the-art methods on all three tasks. Results are shown in Fig. 1. We hope that this paper contributes to accelerating the progress in cross-modal media generation.
@@ 2 Related Work
Diffusion models and Rectified Flow. Starting from Gaussian noise, diffusion [80, 33] and score-based [36, 81] generative models progressively approximate the reverse ODE of a stochastic forward process to generate data. These models have driven significant advances across various domains, particularly in high-fidelity image [17, 35, 62, 5, 54], video [34, 8, 79, 7, 64], and 3D generation [65, 47]. The rapid development of generative models in recent years demonstrates the effectiveness and stability of diffusion models in modeling complex real-world distributions. Recently, rectified flow models [50, 2, 55], such as flow matching, have been proposed to improve the generative process by enabling a transport map between two distributions via ODE. They enable faster training and sampling by avoiding complex probability flow ODEs.
Directly bridging distributions. Flow Matching theoretically allows for arbitrary distributions as the source distribution, which can then be used for direct evolution. Various approaches have been proposed in this direction, such as InterFlow [1], $\alpha$-blending [29], and Schrödinger Bridge [52, 77, 95, 51, 15, 83, 84]. These methods provide important theoretical support for using ODE-based methods to bridge two arbitrary distributions. However, they are still limited to the similar distributions from the same domain, such as image-to-image translation (e.g., faces-to-faces [55, 95] or sketches-to-images [52]). As a step forward, CrossFlow focuses on learning the mapping between data distributions arising from even different modalities.
Text-to-Image generation. Text-to-image generation [68, 59, 74, 73, 69, 10, 22, 14] has witnessed rapid advancements with the advent of diffusion and subsequently flow matching models. This task bridges two complex and important domains: language and vision. To tackle this challenge, existing methods often incorporate text encoders such as LLMs into the diffusion model through additional conditioning mechanisms, with cross-attention being the most widely adopted [22, 64]. However, these approaches increase modeling complexity and require additional parameters. We demonstrate that CrossFlow improves over standard flow matching with better scaling characteristics, and is comparable to prior work, despite a simpler architecture.
@@ 3 Preliminaries
Flow Matching. We consider a generative model that defines a mapping between samples $z_{0}$ from a source distribution $p_{0}$ to samples $z_{1}$ of a target distribution $p_{1}$ via the ordinary differential equation (ODE): $dz_{t}=v_{\theta}(z_{t},t)dt$. Here, $v_{\theta}$ is the velocity parameterized by the weights $\theta$ of a neural network, and $t\in[0,1]$ is the time-step.
Flow Matching [50, 2, 55] defines the forward process using the optimal transport path, i.e.,
@@
where $\sigma_{min}=10^{-5}$. The ground truth velocity can be derived as:
@@
As discussed earlier, flow matching enables evolving a sample $z_{1}$ from an arbitrary source distribution $p_{0}$. But prior work [22, 64] has typically relied on starting from a simple Gaussian noise sample $z_{0}\sim\mathcal{N}(0,1)$, and computing the velocity with additional condition $c$ incorporated through various methods, e.g., cross-attention [22, 64], channel-wise concatenation [27].
@@ 4.2 Training CrossFlow
where $\lambda$ is the weight of KL-divergence loss. Please find more details in Sec. A.1.
@@ 4.3 Classifier-Free Guidance with an Indicator
CFG [32] has become the standard low-temperature sampling method for enhancing multi-modal alignment and improving quality. However, it can only be applied to generation methods that accept an additional conditioning input $c$, since the guidance signal relies on the difference between conditional and unconditional predictions $v_{\theta}(z_{t},c)$ and $v_{\theta}(z_{t})$. Recently, Autoguidance (AG) [38] has been introduced as a method to enhance both conditional and unconditional generation, by leveraging a smaller and less-trained “bad model” as guidance. Nevertheless, it falls short of standard CFG in terms of performance. Moreover, AG requires training a separate bad model, and its performance varies dramatically based on the choice of the bad model. The search space can be restricted by using an under-trained version of the same model, but this affects performance, and also is prohibitive for large models since it requires loading two large models in memory for inference.
We instead aim to support CFG for CrossFlow, which is as accessible and performant as CFG is for standard flow matching. To enable CFG without the presence of an explicit conditioning input $c$, we introduce CFG with an indicator. Specifically, our model is of the form $v_{\theta}(z_{t},1_{c})$, where $1_{c}\in\{0,1\}$ is an indicator to specify conditional vs. unconditional generation. The model evolves from $z_{0}$ to $z_{1}$ when $11_{c}=$ 1, and from $z_{0}$ to $z_{1}^{uc}$ when $01_{c}=$ 0, where $z_{1}^{uc}$ represents any sample from the target distribution $p_{1}$ other than $z_{1}$. During training, we employ two learnable parameters, $g^{c}$ and $g^{uc}$, corresponding to conditional and unconditional generation, respectively. Depending on $1_{c}$, the appropriate learnable parameter is concatenated with the transformer input tokens along sequence dimension. We randomly sample the indicator with an unconditional rate of $10\%$, as per standard practice. The insight behind the CFG indicator is similar to that of standard CFG. In this approach, $v_{\theta}(z_{t},1)$ is trained to map $z_{0}$ to a specific region of the target manifold, while $v_{\theta}(z_{t},0)$ is trained to map $z_{0}$ to the entire target manifold.
@@ 5.1 Text-to-Image Generation
Experimental setup. Comparing T2I models scientifically is challenging due to the diverse datasets used for training, which often include proprietary data, and varying training conditions. In addition, our proposed method represents a new paradigm for utilizing diffusion models, distinct from the previous T2I approaches. Therefore, we primarily compare our model with the widely used “standard flow matching baseline” that starts from noise and leverages text cross-attention. For a fair comparison, we use the exact same codebase, training recipe, dataset, and budget to train both CrossFlow and the baseline. Note that the baseline requires a cross-attention layer after each self-attention layer, whereas our model only relies on self-attention layers, resulting in fewer parameters for the same number of layers. To account for this, we adjust the number of layers to ensure that both models have similar model sizes. For both methods, we use a grid search to find the optimal CFG scale. We also compare CrossFlow with state-of-the-art T2I models to demonstrate that our approach is competitive with those established methods.
Architecture. Our model enables the use of vanilla Transformer [86] with self-attention layers and feed-forward layers. We adopt DiMR [54] as the base architecture for the flow matching models, a variant of Diffusion Transformer (DiT) [62] which replaces the parameter-heavy MLP in adaLN-Zero with a more parameter-efficient Time-Dependent Layer Normalization. For the Text VE, we employ stacked Transformer blocks, followed by a linear layer to project the output into the target shape.
@@
Training details. We use a proprietary dataset with about $350$M image-text pairs to train both CrossFlow and our ablations. Our text encoder is based on CLIP [66] with a fixed sequence length of $77$ text tokens. We use a pre-trained and frozen VAE from LDM [73] to extract image latents. All T2I models are trained using the same settings: an image resolution of $256\times 256$, a batch size of $1024$, a base learning rate of $1\times 10^{-4}$ with $5000$ warm-up steps, and an AdamW optimizer [57] with $\beta_{1}=\beta_{2}=0.9$ and a weight decay of $0.03$. We train our largest model ($0.95$B) on $256\times 256$ for $600$K iterations, then finetune it on $512\times 512$ for an additional $240$K iterations to generate higher-resolution images.
@@
Table 2: Comparison with recent T2I models. For GenEval, we report the overall score here and provide task-specific scores in Sec. B.1. CrossFlow achieves comparable performance with state-of-the-art T2I models by directly evolving text into images.
@@ 5.1.2 State-of-the-art Comparison
Finally, we compare CrossFlow with state-of-the-art text-to-image models and report results in Tab. 2. We achieve a zero-shot FID-30K of $9.63$ on COCO, and a GenEval score of $0.55$, demonstrating performance comparable with the state-of-the-art. Note that our model uses only 630 A100 GPU days for training, whereas other methods like DALL·E 2 [69] typically require thousands of A100 GPU days. These results suggest that CrossFlow is a simple and promising direction for state-of-the-art media generation.
@@ 5.1.3 Arithmetic Operations in Latent Space
Unlike previous diffusion or flow matching models, CrossFlow offers a unique property: arithmetic operations in the input latent space translate to similar operations in the output space. This is made possible since CrossFlow transforms the source space (i.e., the text latent space for T2I) into a regularized continuous space, where a uniform representation shape is shared across all texts. We showcase two examples of this, latent interpolation, and latent arithmetic. For linear interpolation, we use the Text Variational Encoder to generate text latents from two different text inputs, and then interpolate between them to produce images. As shown in Fig. 4, CrossFlow enables visually smooth linear interpolations, even between disparate prompts. Next, we showcase arithmetic operations in Fig. 5, in which we apply addition and subtraction in the text latent space, and find that the resulting images exhibit corresponding semantic modifications to the original image. This demonstrates that CrossFlow formulates meaningful and well-structured semantic paths between the source and the target distributions, providing interesting capabilities and more control compared to standard flow matching approaches.
@@ 5.2 Ablation Study
Text VE loss. We explore reconstruction and contrastive objectives for the encoder loss $L_{Enc}$ when training the text VE. We find that a contrastive loss which promotes semantic understanding yields significantly better performance than reconstruction loss on the input text embeddings. Furthermore, image-text contrastive loss slightly outperforms text-text contrastive loss.
@@ 5.3 CrossFlow for Various Tasks
Image to text (captioning). We first consider the task of image captioning. To achieve this, we train a new Text Variational Encoder on the captioning dataset to extract text latents from text tokens, and a separate text decoder with a reconstruction loss to convert text latents back into tokens. CrossFlow is then trained to map from the image latent space to the text latent space. Following previous work, we use the Karpathy split [37] of COCO dataset [49] for training and testing, and report results in Tab. 4. CrossFlow enables direct evolution from image space to text space for image captioning, achieving state-of-the-art performance.
@@
Table 4: Image captioning on COCO Karpathy split. CrossFlow directly evolves from image to text, achieving comparable performance to state-of-the-art models on image captioning. For a fair comparison, we only consider non-autoregressive methods that are trained without CIDEr optimization.
@@
Low-resolution to high-resolution (super-resolution). We compare CrossFlow with the standard super-resolution method using flow matching, which involves upsampling the low-resolution image and then concatenating it with input noise as conditioning before feeding it into the neural network. In contrast, our method directly evolves the upsampled low-resolution image into a high-resolution image, without additional concatenation conditioning. We also compare against SR3 [75] which uses diffusion models for super-resolution. Following previous work [75, 50], we train and evaluate our model on ImageNet [16] for the task of $64\times 64\rightarrow 256\times 256$ super-resolution, and provide results in Tab. 6. Our method achieves better results compared to the standard flow matching baseline and SR3, indicating that CrossFlow can also effectively evolve between similar distributions while achieving superior performance.
@@
Table 6: Image super-resolution on the ImageNet validation set. Compared with standard SR method with flow matching, our direct mapping method achieves better performance.

### v2

# Flowing from Words to Pixels: A Noise-Free Framework for Cross-Modality Evolution

@@ 1 Introduction
In this work, we present key architecture design solutions for overcoming these challenges: First, we employ a Variational Encoder for encoding the source modality data distribution to the same shape as the target modality, and show that the resulting regularization in the source distribution is essential for generation performance. Secondly, we enable CFG in cross-modal flow matching through the introduction of a binary conditioning indicator during training, and demonstrate the quantitative benefits of this approach compared to alternative CFG methods. We present CrossFlow; a general framework for mapping between two different modalities without the need for any conditioning mechanism or noise distribution. Typically, different cross-modal generation tasks require task-specific architectural and training modifications, but CrossFlow works for different tasks without any such changes.
Using the ubiquitous albeit challenging text-to-image (T2I) generation task as our primary setting, we show the significant result that CrossFlow outperforms commonly used flow matching baselines, given the same training data, model size, and training budget, all without requiring any cross-attention layers. CrossFlow exhibits improved scaling behavior over standard flow matching using cross-attention when scaling training steps or model size, and is also compatible with a variety of Large Language Models (LLMs), including CLIP [70], T5 [71], and Llama3 [20]. Additionally, we demonstrate that since our approach encodes the source distribution into a regularized continuous space with semantic structure, CrossFlow enables exciting new latent arithmetic for the text-to-image task, e.g., $\mathcal{L}$(“A dog with a hat”) + $\mathcal{L}$(“Sunglasses”) – $\mathcal{L}$(“A hat”) creates an image of a dog wearing sunglasses without a hat. Lastly, CrossFlow enables bi-directional mapping between modalities, allowing, for instance, the inversion of text-to-image models to serve as image-to-text (captioning) models.
We demonstrate the general-purpose nature of CrossFlow on various cross-modal/intra-modal tasks: image-to-text (image captioning), image-to-depth (depth estimation), and low-resolution to high-resolution image (super-resolution). CrossFlow achieves comparable or superior performance to various state-of-the-art methods on all three tasks, without requiring task specific architectures. For example, in image captioning, CrossFlow directly projects images into a textual latent space to generate captions, achieving state-of-the-art performance using only a simple text decoder that maps textual latents to discrete tokens. Results are shown in Fig. 1. We hope this paper contributes to accelerating the progress in cross-modal media generation.
@@ 2 Related Work
Diffusion models and rectified flow. Starting from Gaussian noise, diffusion [84, 34] and score-based [37, 85] generative models progressively approximate the reverse ODE of a stochastic forward process to generate data. These models have driven significant advances across various domains, particularly in high-fidelity image [18, 36, 66, 6, 56], video [35, 9, 83, 8, 68], and 3D generation [69, 48, 55, 57]. Recently, rectified flow models [51, 1, 58], such as flow matching, have been proposed to improve the generative process by enabling a transport map between two distributions. They enable faster training and sampling by avoiding complex probability flow ODEs.
Directly bridging distributions. Flow Matching theoretically allows for arbitrary distributions as the source distribution, which can then be used for direct evolution. Various approaches have been proposed in this direction, such as InterFlow [1], $\alpha$-blending [30], data-dependent coupling [3], and Schrödinger Bridge [53, 81, 99, 52, 16, 87, 88]. They provide important theoretical support for using ODE-based methods to bridge two arbitrary distributions. However, they are still limited to similar distributions from the same domain, such as image-to-image translation (e.g., faces-to-faces [58, 99] or sketches-to-images [53]). As a step forward, CrossFlow focuses on learning the mapping between data distributions arising from even different modalities.
Text-to-image generation. Text-to-image generation [72, 63, 78, 77, 73, 11, 101, 23, 15] has rapidly advanced with diffusion and later flow matching models. This task bridges two critical and complex domains: language and vision. Existing methods typically integrate text encoders, such as LLMs, into diffusion models through additional conditioning mechanisms, with cross-attention being the most prevalent [23, 68]. However, these approaches increase model complexity and require extra parameters. We demonstrate that CrossFlow improves over standard flow matching with better scaling characteristics, and is comparable to prior work, despite a simpler architecture.
@@ 3 Preliminaries
Flow Matching. We consider a generative model that defines a mapping between samples $z_{0}$ from a source distribution $p_{0}$ to samples $z_{1}$ of a target distribution $p_{1}$ via the ordinary differential equation (ODE): $dz_{t}=v_{\theta}(z_{t},t)dt$. Here, $v_{\theta}$ is the velocity parameterized by the weights $\theta$ of a neural network, and $t\in[0,1]$ is the time-step. Specifically, Flow Matching [51, 1, 58] defines the forward process as:
@@
and $\sigma_{min}=10^{-5}$. Ground truth velocity is computed as:
@@
As discussed earlier, flow matching enables evolving a sample $z_{1}$ from an arbitrary source distribution $p_{0}$. But prior work [23, 68] typically starts from Gaussian noise $z_{0}\sim\mathcal{N}(0,1)$, and computing the velocity with additional condition $c$ incorporated through various methods, e.g., cross-attention [23, 68], channel-wise concatenation [28].
@@ 4.2 Training CrossFlow
where $\lambda$ is the KL-divergence loss weight. Eq. 4 outlines the general form of the loss function across tasks, where $L_{enc}$ varies by task. Sec. 4.4 discusses text-to-image generation and choices for $L_{Enc}$. More details in Sec. A.1.
@@ 4.3 Classifier-Free Guidance with an Indicator
CFG [33] has become the standard low-temperature sampling method for enhancing multi-modal alignment and improving quality. However, it can only be applied to generation methods that accept an additional conditioning input $c$, since the guidance signal relies on the difference between conditional and unconditional predictions $v_{\theta}(z_{t},c)$ and $v_{\theta}(z_{t})$. Recently, Autoguidance (AG) [39] has been introduced as a method to improve both conditional and unconditional generation, by guiding with a smaller, less-trained ‘bad model’. However, it underperforms compared to standard CFG. AG also requires training a separate bad model, and its performance varies dramatically based on the choice of the bad model. While using an under-trained version of the same model narrows the search space, it affects performance and is impractical for large models due to the need to load two models during inference.
We instead aim to support CFG for CrossFlow, which is as accessible and performant as CFG is for standard flow matching. To enable CFG without the presence of an explicit conditioning input $c$, we introduce CFG with indicator. Specifically, our model is of the form $v_{\theta}(z_{t},1_{c})$, where $1_{c}\in\{0,1\}$ is an indicator to specify conditional vs. unconditional generation. The model evolves from $z_{0}$ to $z_{1}$ when $11_{c}=$ 1, and from $z_{0}$ to $z_{1}^{uc}$ when $01_{c}=$ 0, where $z_{1}^{uc}$ represents any sample from the target distribution $p_{1}$ other than $z_{1}$. During training, we employ two learnable parameters, $g^{c}$ and $g^{uc}$, corresponding to conditional and unconditional generation, respectively. Depending on $1_{c}$, the appropriate learnable parameter is concatenated with the transformer input tokens along sequence dimension. We randomly sample the indicator with an unconditional rate of $10\%$, as per standard practice. The insight behind the CFG indicator is similar to that of standard CFG. In this approach, $v_{\theta}(z_{t},1)$ is trained to map $z_{0}$ to a specific region of the target manifold, while $v_{\theta}(z_{t},0)$ is trained to map $z_{0}$ to the entire target manifold to generate arbitrary unrelated images.
@@ 5.1 Text-to-Image Generation
Experimental setup. Scientifically comparing T2I models is challenging due to diverse training datasets, often including proprietary data, and varying training conditions. In addition, our method represents a new paradigm for utilizing diffusion models, distinct from previous T2I approaches. Therefore, we primarily compare our model with the widely used “standard flow matching baseline" that starts from noise and uses text cross-attention. For fairness, both CrossFlow and the baseline share the same codebase, training recipe, dataset, and budget. Unlike the baseline, which requires cross-attention after each self-attention layer, our model only relies on self-attention, reducing parameters per layer. To account for this, we adjust the number of layers to match model sizes. For both methods, we use a grid search to find the optimal CFG scale. We also compare CrossFlow with state-of-the-art T2I models to demonstrate that our approach is competitive with those established methods.
Architecture. Our model enables the use of vanilla Transformer [90] with self-attention layers and feed-forward layers. We use DiMR [56] as the flow matching backbone, a variant of Diffusion Transformer (DiT) [66] which replaces the parameter-heavy MLP in adaLN-Zero with a lightweight Time-Dependent Layer Normalization. For the Text VE, we apply stacked Transformer blocks, followed by a linear layer to project the output into the target shape.
@@
Training details. We use a proprietary dataset with about $350$M image-text pairs to train both CrossFlow and our ablations. Our text encoder is based on CLIP [70] with a fixed sequence length of $77$ text tokens. We use a pre-trained and frozen VAE from LDM [77] to extract image latents. Logit-normal sampling [23] is used to bias the training timesteps. All T2I models are trained using the same settings: an image resolution of $256\times 256$, a batch size of $1024$, a base learning rate of $1\times 10^{-4}$ with $5000$ warm-up steps, and an AdamW optimizer [60] with $\beta_{1}=\beta_{2}=0.9$ and a weight decay of $0.03$, and a KL loss weight of $\lambda=1\times 10^{-4}$. We train our largest model ($0.95$B) on $256\times 256$ for $600$K iterations, then finetune it on $512\times 512$ for an additional $240$K iterations for higher resolution generation.
@@
CrossFlow (Sin-Cos) 0.95B 8.95 0.57
Table 2: Comparison to recent T2I models. For GenEval, we report overall scores here and task-specific results in Sec. B.1. CrossFlow achieves comparable results to state-of-the-art models by evolving text directly into images. CrossFlow (Sin-Cos) replaces simple linear flow matching with sin-cos matching [2].
@@ 5.1.2 State-of-the-art Comparison
Finally, we compare CrossFlow with state-of-the-art text-to-image models and report results in Tab. 2. We additionally explore sin-cos matching [2] and find it improves over vanilla linear flow matching. We achieve a zero-shot FID-30K of $8.95$ on COCO, and a GenEval score of $0.57$, demonstrating performance comparable with the state-of-the-art. Note that our model uses only 630 A100 GPU-days for training, whereas other methods like DALL·E 2 [73] typically require thousands of A100 GPU days. These results suggest that CrossFlow is a simple and promising direction for state-of-the-art media generation.
@@ 5.1.3 Arithmetic Operations in Latent Space
Unlike previous diffusion or flow matching models, CrossFlow offers a unique property: arithmetic operations in the input latent space translate to similar operations in the output space. This is made possible since CrossFlow transforms the source space (i.e., the text latent space for T2I) into a regularized continuous space, where a uniform representation shape is shared across all texts. We showcase two examples of this, latent interpolation, and latent arithmetic. For latent interpolation, we use the Text Variational Encoder to generate text latents from two different text inputs, and then interpolate between them to produce images. As shown in Fig. 4, CrossFlow enables visually smooth linear interpolations, even between disparate prompts. Next, we showcase arithmetic operations in Fig. 5, in which we apply addition and subtraction in the text latent space, and find that the resulting images exhibit corresponding semantic modifications to the original image. This shows that CrossFlow formulates meaningful and well-structured semantic paths between the source and target distributions, providing additional capabilities and more control over standard flow matching approaches. See Sec. B.2 for further details.
@@ 5.2 Ablation Study
Text VE loss. We explore reconstruction and contrastive objectives for the encoder loss $L_{Enc}$ when training the text VE. We find that contrastive loss, which enhances semantic understanding, significantly outperforms reconstruction loss on input text embeddings. Moreover, image-text contrastive loss slightly surpasses text-text contrastive loss.
@@ 5.3 CrossFlow for Various Tasks
Image to text (captioning). We first consider the task of image captioning. To achieve this, we train a new Text Variational Encoder on the captioning dataset to extract text latents from text tokens, and a separate text decoder with a reconstruction loss to convert text latents back into tokens. CrossFlow is then trained to map from the image latent space to the text latent space. Following previous work, we use the Karpathy split [38] of COCO dataset [50] for training and testing. In addition, we can also leverage the bi-directional flow property, and simply fine-tune our text-to-image CrossFlow model on COCO and use its inversion for captioning. We report results in Tab. 4. CrossFlow enables direct evolution from image space to text space for image captioning, achieving state-of-the-art performance.
@@
Table 4: Image captioning on COCO Karpathy split. CrossFlow directly evolves from image to text, achieving comparable performance to state-of-the-art models on image captioning. For a fair comparison, we consider non-autoregressive methods that are trained without CIDEr optimization. CrossFlow-T2I achieves captioning by simply inverting our text-to-image CrossFlow model.
@@
Low-resolution to high-resolution (super-resolution). We compare CrossFlow with the standard flow-matching super-resolution method, which upsamples the low-resolution image, concatenates it with input noise as conditioning, and then processes it through the neural network. In contrast, we directly evolve the upsampled low-resolution image into a high-resolution image, without additional concatenation conditioning. We also compare against SR3 [79] which uses diffusion models for super-resolution. Following previous work [79, 51], we train and evaluate our model on ImageNet [17] for $64\times 64\rightarrow 256\times 256$ super-resolution, and provide results in Tab. 6. Our method achieves better results compared to the standard flow matching and SR3, indicating that CrossFlow can also effectively evolve between similar distributions while achieving superior performance.
@@
Table 6: Image super-resolution on the ImageNet validation set. Our direct mapping method achieves better performance.